OpenStudy (anonymous):

find the exact value of sin(5pi/12). i really need to check my work because something went wrong

OpenStudy (paki):

0.022

OpenStudy (anonymous):

i need to find the exact value though

OpenStudy (paki):

yeah its the exact one.... 0.02283274 :-)

OpenStudy (anonymous):

as in I need to apply trig addition identities

OpenStudy (ranga):

5*pi/12 radians = 5*pi/12 * 180 / pi degrees = 75 degrees. sin(75) = sin(45 + 30) Use trig addition identity to expand sin(45 + 30).

OpenStudy (ranga):

sin(A+B) = sin(A)cos(B) + cos(A)sin(B)

OpenStudy (anonymous):

yes i used that in radians. but i apparently got the wrong answer

OpenStudy (anonymous):